1) a) Mitra Co. is trying to determine their costs of preferred and common. Its common stock sells for $50 per share and will pay a $6 dividend which is expected to grow at a constant 5% rate. Its preferred stock sells for $22.50 per share and pays $1.80 in dividends.

Calculate a firm's required rates of return for both of its equity components. (6 marks)

b) What accounts for the difference in returns, given that these are both forms of equity? (2 marks)

PLEASE SHOW ALL WORK AND CALCULATIONS
Answer #1

a) Cost of Common Equity can be calculated by using the constant growth dividend discount model, which is mathematically represented as:

V_0 = \frac{D_1}{r - g}

where D1 is dividend expected = $6 in question, r is the cost of equity or the required return and g is the growth rate

50 = \frac{6}{r - 0.05}

r = (6/50) + 0.05 = 17%

Cost of preferred equity = Annual dividend/Current Price of Share = 1.80/22.50 = 8%

b) Difference in returns of two forms of equity signifies different levels of risk that are attached with these forms. Preferred stocks are ranked higher in terms of claims on assets than common stock. This implies, in event of bankruptcy or liquidation, preferred shareholders will be paid first and than common stock holders will be paid.

This would mean that the risk attached to common equity would be higher. By basic finance principle, higher risk would mean higher expected return and hence, common equity having higher risk, has higher required rate or cost attached to it.